SpaceRef

SpaceRef


NASA HST Advanced Camera for Surveys (ACS) Status Update

Status Report From: Goddard Space Flight Center
Posted: Thursday, June 29, 2006

image

On Thursday afternoon, June 29, the HST's senior managers attended the Flight Readiness Review (FRR) of the activities required to install flight software into, and transition the ACS to its Side 2 (redundant) electronics. The review was also attended by Dr. Jennifer Wiseman (NASA HQ) and the ST ScI Director, Dr. Matt Mountain.

The first portion of the meeting was devoted to a status report from the Anomaly Review Board investigating the suspension of ACS science operations on June 19. The report summarized the current findings of the ARB, and discussed its hypothesis that the June 19 anomaly was triggered by a loss of the +15 Volt source provided by ACS's Low Voltage Power Supply 1. The ARB chairman, Roger Chiei then reported the ARB's unanimous conclusion that it is safe use the Side 2 electronics to recover ACS to science operations. The ARB also recommended that the HST Program proceed accordingly. It added the caveat that some specific fault scenarios would preclude success of the transition to Side 2 without posing any risk of additional damage to the instrument. A successful recovery will rule out these specific faults. Finally, the ARB recommended that there be a temporary moratorium on use of the ACS's Solar Blind Channel (SBC) while the ARB completes additional analysis of the ACS electronics.

After discussion, HST's management adopted all of the ARB's recommendations and turned to reviewing readiness to proceed with the recovery. The FRR itself was completed about 3:30 p.m. At it's conclusion, HST's Operations Project directed initiation of ACS recovery operations. Commanding commenced this evening around 5 PM, and activities are expected to complete by 4 PM, EDT on Friday, June 30. The first major activity is the loading of ACS flight software into the Side 2 EEPROM. The last activities will include a brief checkout of ACS Side 2 functionality followed by placement of ACS into its Safe Mode on Friday afternoon. ACS will be then be returned to Operate Mode on Sunday evening. Measurements of dark and bias counts will be made on July 3 and 4, and ACS science observations will resume on Tuesday evening, July 4.

Additional status updates will be issued after the planned activation of the ACS CCD Electronics Boxes (CEBs) on Friday afternoon, after Sunday evening's recovery of ACS from SAFE to OPERATE and after evaluation of the data from the initial exposures made with ACS's CCD detectors (i.e., around the middle of the coming week.

Dr. J. Keith Kalinowski
Deputy Manager, Hubble Space Telescope Operations Project
NASA/GSFC Code 441
Building 7, Room 263
Greenbelt, MD 20771

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.